### v2

# Local Order Average-Atom Interatomic Potentials

Abstract
This article describes an extension to the effective Average Atom (AA) method for random alloys to account for local ordering (short-range order) effects by utilizing information from partial radial distribution functions. The new Local-Order Average Atom (LOAA) method is rigorously derived based on statistical mechanics arguments and validated for non-stoichiometric binary 2D hexagonal crystals and 3D FeNiCr and NiAl alloys whose ground state is obtained through Monte Carlo sampling. Material properties for these alloys, and phase transformations for the NiAl system, computed from static and dynamic atomistic simulations using standard interatomic potentials (IPs) exhibit a strong dependence on local ordering that is captured by simulations with effective LOAA IPs, but not the original AA method. The advantage of LOAA is that it requires smaller system sizes to achieve statistically converged results and therefore enables the simulation of complex materials, such as high-entropy alloys, at a fraction of the computational cost of standard IPs.
preprint: APS/123-QED
I Introduction
The average-atom (AA) approach to modeling random alloys is based on an effective medium approximation in which the species of atoms are taken to be independent and consistent with their concentrations in a crystalline alloy. In this method, each atom is identical and typeless, with an effective atomic interatomic potential (IP) in which interactions between atoms are computed as an average over interactions that account for species identities weighted by their probability of occurrence assuming an ideal random alloy. The AA approach was originally introduced by Smith and Was [smith:was:1989], Ackland and Vitek [ackland:vitek:2009], and Najafabadi et al. [najafabadi:wang:1991], and revived by Curtin and coworkers [varvenne:luque:2016, nohring:curtin:2016]. The method has primarily been applied to alloy systems to study the effect of species concentrations on material properties [jian:xie:2020, zhao:xiong:2021, cheng:yuan:2021, seoane:farkas:2022, liang:goodelman:2023, fang:li:2024]. A recent application area for AA is the class of high entropy-alloys (HEAs) [George2019-pl] or multi-principal element alloys [Yeh2004-dy, Shittu2024-wb]. These are alloys that typically contain five or more species of approximately equal concentrations and which are of interest due to their high strength and high fracture toughness relative to traditional alloys [yeh:chen:2004, gludovatz:hohenwarter:2014, zhang:liu:2024]. Such systems require large simulation cells in order to correctly capture species fractional occupancy and hence benefit from AA, which can provide converged results with much smaller system sizes.
While the AA approach provides an excellent model for the ideal case of random alloys, in practice all materials exhibit some level of “local ordering” (LO) (also called “short-range order”), i.e. a preferential arrangement making certain species combinations at certain distances more or less favorable than the ideal random case [Sheriff2024-jk, Islam2025]. Bragg and Williams [bragg:williams:1934] first noticed the effects of temperature on atomic arrangements in the 1930s. Since then, it has become clear that LO exists in alloys, and that it can have important effects on material properties [mcmanus:1965]. For example, in the brittle-to-ductile transition, specific LO arrangements of atoms can directly result in increased stiffness while maintaining ductility [Wang2023-qy]. Liquid alloys are another example where LO becomes important. During the cooling process of an alloy, chemical short-range order plays a critical role in both the nucleation mechanism and the specific phases and microstructures formed [Ruppersberg1975, Liang2020-ql, Greer2016, Simonet1998-rm]. More generally, ordering in liquid alloys affects the electrical, thermodynamic, and structural properties [Saboungi1990]. LO is also very important for HEAs, as it can dramatically change a material’s strength, hardness, ductility, and defect evolution [zhang:zhao:2020, chen:aitken:2021, zhao:2021]. Understanding LO in HEAs and other alloys is therefore important for predicting their response to changes in stress and temperature for real-world applications.
Here we extend the AA formalism to explicitly account for LO effects by including information from partial radial distribution functions (RDFs) in the definition of the effective IP. The new local-order average-atom (LOAA) method is as efficient as AA, but correctly captures the effect of LO on alloy properties. In LOAA, the effective interaction between atoms is weighted based on the probabilities of observing pairs of species at a given distance as described by the partial RDFs, as opposed to ideal random alloy probabilities. We validate the approach by studying the elastic properties of a two-dimensional (2D) binary hexagonal crystal with Lennard-Jones (LJ) interactions, a three-dimensional (3D) Fe(1-x)/2Ni(1-x)/2Crx alloy with embedded-atom method (EAM) interactions, and martensitic phase transformations in off-stochiometric NiAl with EAM interactions. We compare the accuracy of LOAA with AA, and with “true species” (TS) computations, calculations using standard IPs with the actual atomic species identities, averaged over many realizations of the system. Overall we find LOAA to provide a more accurate estimate for the TS case than AA with no increase in computational cost except for the initial one-time generation of the partial RDFs. We observe that while the AA method can only accurately predict properties such as lattice constants, energies, and elastic constants for an ideal random alloy, the LOAA formulation correctly predicts these properties with increasing LO.
The remaining sections of this paper are organized as follows. Section II introduces the basic definitions and derivation of the LOAA method. Section III includes results for a 2D binary system where the effect of LO is explored by tuning the LJ parameters, and 3D alloy systems (FeNiCr and NiAl) where the effect of LO on mechanical properties and phase transformations is studied. Section IV concludes with a summary and a discussion of possible extensions to the LOAA method. The appendices contain additional derivations and results referred to in the main sections of the paper.
II Local Order Average-Atom Potentials
II.1 Preliminaries
Consider an infinite multilattice crystal (i.e. a lattice with a basis) with species $A$, $B$, …appearing in concentrations $c_{A}$, $c_{B}$, …, such that
$\lim_{N\to\infty}\frac{1}{N}\sum_{i=1}^{N}s_{i}^{X}=c_{X},\quad x=A,B,\dots,$ (1)
where $N$ is the number of atoms in the crystal, and
$s_{i}^{X}=\begin{cases}1&\text{if atom $i$ is of species $X$},\\ 0&\text{otherwise}.\end{cases}$ (2)
We define the ensemble average over all possible species arrangements $\mathcal{S}$ satisfying the constraints in Eq. (1), for a given geometric configuration $\mathcal{C}$ specifying the positions of the atoms, as
$\langle f(i,j,\dots);\mathcal{C}\rangle_{\mathcal{S}}=\lim_{M\to\infty}\frac{1}{M}\sum_{m=1}^{M}f(i,j,\dots;\mathcal{S}_{m}),$ (3)
where $f(i,j,\dots;\mathcal{S}_{m})$ is some function related to atoms $i,j,\dots$ that may depend on their species, $\mathcal{S}_{m}$ is a particular species arrangement (realization) on configuration $\mathcal{C}$, and $M$ is the size of the ensemble taken to be infinite since the system size is infinite.
Eq. (1) is the concentration constraint for a single species arrangement for the specified configuration $\mathcal{C}$. This constraint implies that the probability $p(s_{i}^{X}=1\,|\,\mathcal{C})$ that atom $i$ is of type $X$ given geometric configuration $\mathcal{C}$ is $c_{X}$. [^1] This is evident if one considers drawing atom $i$ out of the pool of all possible realizations where the fraction of atoms of species $X$ is $c_{X}$. The expectation (average value) of $s_{i}^{X}$ follows by applying Eq. (3):
$\langle s_{i}^{X};\mathcal{C}\rangle_{\mathcal{S}}=1\times c_{X}+0\times(1-c_{X})=c_{X}.$ (4)
II.2 Pairwise Interactions
The energy of a particular species arrangement on the crystal assuming pairwise interactions is
$E=\frac{1}{2}\sum_{i}\sum_{j\neq i}\sum_{X}\sum_{Y}V_{ij}^{XY}s_{i}^{X}s_{j}^{Y},$ (5)
where $i$ and $j$ run over all atoms, and $X$ and $Y$ over all species, and $V_{ij}^{XY}$ is the energy in a bond connecting pair $ij$ of species $XY$,
$V_{ij}^{XY}=V_{XY}(r_{ij}),$ (6)
where $V_{XY}(r)$ is the pair potential function for species $X$ and $Y$, and $r_{ij}$ is the distance between atoms $i$ and $j$.
First, we consider the random alloy case where atomic species are randomly assigned to atoms consistent with a specific set of concentrations. The average value (expectation) of the energy follows by applying the ensemble average in Eq. (3):
$\langle E;\mathcal{C}\rangle_{\mathcal{S}}$ $=\left\langle\frac{1}{2}\sum_{i}\sum_{j\neq i}\sum_{X}\sum_{Y}V_{ij}^{XY}s_{i}^{X}s_{j}^{Y};\mathcal{C}\right\rangle_{\mathcal{S}}$
$=\frac{1}{2}\sum_{i}\sum_{j\neq i}\sum_{X}\sum_{Y}V_{ij}^{XY}\left\langle s_{i}^{X}s_{j}^{Y};\mathcal{C}\right\rangle_{\mathcal{S}}$
$=\frac{1}{2}\sum_{i}\sum_{j\neq i}\sum_{X}\sum_{Y}V_{ij}^{XY}\left\langle s_{i}^{X};\mathcal{C}\right\rangle_{\mathcal{S}}\left\langle s_{j}^{Y};\mathcal{C}\right\rangle_{\mathcal{S}}$
$=\frac{1}{2}\sum_{i}\sum_{j\neq i}\sum_{X}\sum_{Y}V_{ij}^{XY}c_{X}c_{Y},$ (7)
where in passing from the second to the third line we assume that the variables $s_{i}^{X}$ are independent random variables, i.e. that the species of different atoms are independent, [^2] A necessary (but not sufficient) condition for this is that the system is infinite. As a demonstration of this requirement, consider a crystal with just two atoms ($N=2$) with species AA and $B$, so that the concentrations are $c_{A}=c_{B}=0.5$. According to the reasoning in Eq. (7), the probability that both atoms are of species AA is $c_{X}c_{Y}=0.25$, but this is impossible since there is only one atom of species $A$. The discrepancy is because in a finite crystal occupation of sites is coupled. In this case, if one atom is of type $A$, the other can only be of type $B$, so the probability of it being of type AA is zero not $c_{A}$. Even in an infinite system, it is possible that there is a correlation between atoms due to LO of species. and in passing from the third to fourth line, Eq. (4) is used.
In Eq. (7), we recover the AA formalism in which the energy of the system is expressed in terms of an effective pair potential between typeless atoms (i.e. with no specified species):
$\langle E\rangle_{\mathcal{S}}=\frac{1}{2}\sum_{i}\sum_{j\neq i}V_{\rm eff}(r_{ij}),$ (8)
where
$V_{\rm eff}(r)=\sum_{X}\sum_{Y}V_{XY}(r)c_{X}c_{Y}.$ (9)
In general, Eq. (9) has to be evaluated explicitly, making this approach somewhat more expensive computationally compared to the underlying pair potentials. [^3] For a system containing $n$ species, there are $n(n+1)/2$ function evaluations (accounting for symmetry). Pair potentials with a finite cut-off scale linearly with the number of atoms, $N$. Thus the AA formulation scales as $O(n^{2}N)$. However, as it is more general (and more efficient) to store pair potentials on a discrete grid that is interpolated on-the-fly at runtime, it is usually possible to pre-compute $V_{\rm eff}$ and therefore recover the same speed and scaling of the original pair potential model. The additional benefit of the AA approach, however, is that better statistics are obtained than in a calculation of a single realization of a finite-sized random alloy. One is effectively calculating the average behavior over all such possible realizations.
In the local-order average atom (LOAA) method introduced here, we relax the assumption in Eq. (7) that the variables $s_{i}^{X}$ are independent. In general,
$\left\langle s_{i}^{X}s_{j}^{Y};\mathcal{C}\right\rangle_{\mathcal{S}}$ $=(1\times 1)\,p(s_{i}^{X}=1,s_{j}^{Y}=1\,|\,\mathcal{C})$
$+(1\times 0)\,p(s_{i}^{X}=1,s_{j}^{Y}=0\,|\,\mathcal{C})$
$+(0\times 1)\,p(s_{i}^{X}=0,s_{j}^{Y}=1\,|\,\mathcal{C})$
$+(0\times 0)\,p(s_{i}^{X}=0,s_{j}^{Y}=0\,|\,\mathcal{C})$
$=p(s_{i}^{X}=1,s_{j}^{Y}=1\,|\,\mathcal{C}),$ (10)
where $p(s_{i}^{X}=1,s_{j}^{Y}=1\,|\,\mathcal{C})$ is the probability that atom $i$ is of species $X$ and atom $j$ is of species $Y$ given the geometric configuration $\mathcal{C}$. In special case that the species assignments are independent, then
$p(s_{i}^{X}=1,s_{j}^{Y}=1\,|\,\mathcal{C})=c_{X}c_{Y},$ (11)
so that from Eq. (10),
$\left\langle s_{i}^{X}s_{j}^{Y};\mathcal{C}\right\rangle_{\mathcal{S}}=c_{X}c_{Y},$ (12)
and Eq. (7) is recovered. Otherwise the probability in Eq. (10) must be known. To proceed, based on objectivity requirements, we postulate that the probability does not depend on the identities of the atoms, but only on their species and on the distance between them. [^4] More generally, the probability can depend not just on the two atoms forming the bond, but on distances to and between surrounding atoms and their species. However, given the pair potential interactions considered in this section, we postulate a simple pairwise form. We assume the following functional form:
$p(s_{i}^{X}=1,s_{j}^{Y}=1\,|\,\mathcal{C})=c_{X}c_{Y}\Phi_{XY}(r_{ij};\mathcal{C}).$ (13)
The function $\Phi_{XY}(r;\mathcal{C})$ characterizes the probability of finding a bond of length $r$ connecting species $X$ and $Y$ in geometric configuration $\mathcal{C}$ relative to a random alloy where the species are independent. $\Phi_{XY}(r;\mathcal{C})$ has the following properties:
  1. 1.
    The dependence of $\Phi_{XY}$ on $\mathcal{C}$ affects the functional form. For example if $\mathcal{C}$ is a perfectly crystalline configuration, then atoms are only found at an infinite set of discrete distances (the neighbor shell radii). Therefore $\Phi_{XY}(r;\mathcal{C})=0$ for all distances not in $\mathcal{C}$.
  2. 2.
    $\Phi_{XY}(0;\mathcal{C})=0$. This indicates that two atoms cannot occupy the same position.
  3. 3.
    For a system with local order, i.e. short-range order, $\lim_{r\to\infty}\Phi_{XY}(r;\mathcal{C})=\delta(r;\mathcal{C})$, where $\delta(r;\mathcal{C})=1$ for all distances present in $\mathcal{C}$ and zero otherwise. This indicates that for atoms in $\mathcal{C}$ that are far apart, the probability that they are of species $X$ and $Y$ reverts to the independent case. [^5] We expect this limit to decay quickly with a dependence on the cutoff radius $r_{\rm cut}^{XY}$ of the pair potential $V_{XY}(r)$.
  4. 4.
    $\Phi_{XY}(r;\mathcal{C})=\Phi_{YX}(r;\mathcal{C})$. Symmetry with respect to atomic species is a consequence of objectivity.
Under conditions of thermodynamic equilibrium, the functional form of $\Phi_{XY}(r;\mathcal{C})$ reflects the nature of bonding in the material and the thermodynamic constraints imposed by a statistical mechanics ensemble. We relate the probability of finding atoms of a given species at a given separation to the partial RDFs of the system (see Appendix A for the derivation). The result is
$p(s_{i}^{X}=1,s_{j}^{Y}=1\,|\,\mathcal{C})=\frac{c_{X}c_{Y}g_{XY}(r_{ij})}{\sum_{A}\sum_{B}c_{A}c_{B}g_{AB}(r_{ij})},$ (14)
where $g_{XY}(r_{ij})$ is the partial RDF between species $X$ and $Y$. In general, the RDF is defined in terms of a phase average evaluated over the relevant ensemble, [^6] We provide a general definition here for arbitrary conditions, although in practice we will only consider cases where the RDF is evaluated under static conditions at 0 K. such as the canonical ensemble [tuckerman:2010]. This phase average is evaluated for one specific species arrangement satisfying the concentration constraint in Eq. (1). It can be obtained from a molecular simulation or experiment for a given system at a given temperature $T$. Thus the left-hand side of Eq. (14) is the probability over all species arrangements for a single geometric configuration $\mathcal{C}$, whereas the right-hand side involves all geometric configurations with a single species arrangement $\mathcal{S}$. [^7] This means that each atom is assigned a species, such that Eq. (1) is satisfied, and then the phase average is over all possible arrangements of these atoms. The requirement that this equality holds is of the nature of an ergodic hypothesis. Since $\mathcal{C}$ is a system of infinite extent (see Section II.1), it is assumed that the probability over all species arrangements for a single geometric configuration snapshot at finite temperature (left-hand side of Eq. (14)) is equal to the probability over all geometric configurations for a specified set of species (right-hand side of Eq. (14)).
Comparing Eq. (14) with Eq. (13), we have
$\Phi_{XY}(r)=\frac{g_{XY}(r)}{\sum_{A}\sum_{B}c_{A}c_{B}g_{AB}(r)},$ (15)
where we have dropped the explicit dependence of $\Phi_{XY}(r)$ on the geometric configuration $\mathcal{C}$ since it is now assumed to apply to all configurations (positions and species) consistent with the canonical ensemble that defines the RDF. For the case where the species are independent, all partial RDFs are equal (see Appendix A) and Eq. (15) gives $\Phi_{XY}(r)=1$ as expected (see Eq. (56)).
The derivation leading to Eq. (14) applies to the configuration at which the RDF is evaluated, which is typically a reference equilibrium state. To define an effective LOAA potential, we therefore apply Eq. (14) in the reference configuration. Thus referring to Eqns. (7) and (10), we have
$V_{\rm eff}(r,R)=\sum_{X}\sum_{Y}V_{XY}(r)G_{XY}(R),$ (16)
where $r$ and $R$ are the distance between a pair of atoms in the deformed and reference configurations, respectively, and
$G_{XY}(R)=\frac{c_{X}c_{Y}g_{XY}(R)}{\sum_{A}\sum_{B}c_{A}c_{B}g_{AB}(R)}.$ (17)
Thus if the partial RDFs for an alloy are known (either from experiment or from computation as shown in Section III), it is possible to construct an effective potential that accounts for species ordering. We make the following observations regarding $V_{\rm eff}(r,R)$:
  1. 1.
    The effective potential remains pairwise. It is a combination of the species-dependent pair potentials weighted by the species probabilities determined in the reference state.
  2. 2.
    The effective potential can be used in arbitrary deformed states corresponding to the same reference configuration. [^8] Consider a reference configuration $\mathcal{C}$ that is mapped by some deformation to a new deformed configuration $\mathcal{C}^{\prime}$. Given two atoms $i$ and $j$, the probability that $i$ is of species $X$ and $j$ is of species $Y$ is the same in $\mathcal{C}$ and $\mathcal{C}^{\prime}$, i.e. $\left\langle s_{i}^{X}s_{j}^{Y};\mathcal{C}\right\rangle_{\mathcal{S}}=\left\langle s_{i}^{X}s_{j}^{Y};\mathcal{C}^{\prime}\right\rangle_{\mathcal{S}}$. This is because for any arrangement of species sampled on $\mathcal{C}$, atoms $i$ and $j$ have the same species in $\mathcal{C}^{\prime}$ since the applied deformation does not change species identities. This statement applies at sufficiently low temperatures where species arrangements do not re-equilibrate through diffusion over the time scale being simulated. In other words, the species probabilities between pairs of atoms prior to deformation are “frozen in.”
  3. 3.
    The effective potential is temperature-dependent since the RDFs are obtained within the canonical ensemble at a temperature $T$, so that $G_{XY}=G_{XY}(R;T)$. However, the temperature dependence of the radial distribution is rather weak (see for example the numerical results in [tuckerman:2010]), so $V_{\rm eff}$ can be used for a range of “close” temperatures.
  4. 4.
    Unlike a standard IP, $V_{\rm eff}(r,R)$ requires knowledge of both deformed and reference coordinates:
    $\langle E\rangle_{\mathcal{S}}=\frac{1}{2}\sum_{i}\sum_{j\neq i}V_{\rm eff}(r_{ij},R_{ij}).$ (18)
    In a simulation, this requires extra storage for the atom reference coordinates (or displacement vectors) and modifications to molecular dynamics codes.
  5. 5.
    Computationally, the LOAA approach is not significantly slower than using the AA effective potential for the independent species case defined in Eq. (9), since $G_{XY}(R)$ can be precomputed. [^9] In the following, we will consider cases where the reference configuration is a perfect crystal. In those cases, LOAA can be made virtually as efficient as the underlying pair potential by precomputing a discrete set of $V_{\rm eff}$ functions, one for each of the discrete pair distances in the lattice up to a conservative neighbor distance allowing for atomic motion. Then during a run, the computation only requires a pair to be mapped to its reference configuration, the value of $R$ computed, and the appropriate $V_{\rm eff}$ function accessed. We note that if the only objective is to determine ground state and the partial RDFs are not known from experiments, then LOAA does not provide a computational benefit over brute force TS simulations since the RDF needs to be obtained. However once obtained, the effective LOAA potential can be used to perform a variety of computations on smaller systems efficiently.
II.3 EAM-style Embedding Function
Next consider an energy with an EAM-style nonlinear embedding term,
$E=\sum_{i}\sum_{X}U^{X}(\rho_{i})s_{i}^{X},$ (19)
where $U^{X}$ is the embedding function for an atom of type $X$ and $\rho_{i}$ is the charge density at the position of atom $i$ given by
$\rho_{i}=\sum_{j\neq i}\sum_{Y}\rho_{ij}^{Y}s_{j}^{Y},$ (20)
in which $\rho_{ij}^{Y}$ is the contribution from atom $j$ of species $Y$ to the charge density at atom $i$. The average energy is
$\langle E\rangle_{\mathcal{S}}$ $=\left\langle\sum_{i}\sum_{X}U^{X}(\rho_{i})s_{i}^{X}\right\rangle_{\mathcal{S}}$
$=\sum_{i}\sum_{X}\left\langle U^{X}(\rho_{i})s_{i}^{X}\right\rangle_{\mathcal{S}}$
$=\sum_{i}\sum_{X}\left\langle U^{X}(\rho_{i})\right\rangle_{\mathcal{S}}\left\langle s_{i}^{X}\right\rangle_{\mathcal{S}}$
$=\sum_{i}\sum_{X}\left\langle U^{X}(\rho_{i})\right\rangle_{\mathcal{S}}c_{X}.$ (21)
The passage for the second to third line follows since the terms $U^{X}(\rho_{i})$ and $s_{i}^{X}$ are independent. This is because the density $\rho_{i}$, given in Eq. (20), depends only on the species of the atoms surrounding atom $i$ and not on its species as indicated by $s^{X}_{i}$.
We note that $\left\langle U^{X}(\rho_{i})\right\rangle_{\mathcal{S}}$ does not depend on LO. This is because, as explained above, $\rho_{i}$ is a sum over the electron density contributions of neighboring atoms, each of which depends only on the species of the contributing atom. Therefore radial distributions between atom species do not enter into the embedding term. Thus we can make the interesting observation that LO enters only through the pairwise interactions in EAM functional forms.
An approximate expression for $\left\langle U^{X}(\rho_{i})\right\rangle_{\mathcal{S}}$ can be obtained by expanding the embedding energy about the average charge density $\bar{\rho}_{i}$,
$U^{X}(\rho_{i})=U^{X}(\bar{\rho}_{i})+\left.\frac{dU^{X}}{d\rho}\right|_{\bar{\rho}_{i}}(\rho_{i}-\bar{\rho}_{i})+\frac{1}{2}\left.\frac{d^{2}U^{X}}{d\rho^{2}}\right|_{\bar{\rho}_{i}}(\rho_{i}-\bar{\rho}_{i})^{2}+\dots,$ (22)
where
$\bar{\rho}_{i}=\langle\rho_{i}\rangle_{\mathcal{S}}=\left\langle\sum_{j\neq i}\sum_{Y}\rho_{ij}^{Y}s_{j}^{Y}\right\rangle_{\mathcal{S}}=\sum_{j\neq i}\sum_{Y}\rho_{ij}^{Y}\langle s_{j}^{Y}\rangle_{\mathcal{S}}=\sum_{j\neq i}\sum_{Y}\rho_{ij}^{Y}c_{Y}.$ (23)
Using Eq. (22) and retaining terms to second order, the average embedding energy follows as
$\left\langle U^{X}(\rho_{i})\right\rangle_{\mathcal{S}}$ $\approx\left\langle U^{X}(\bar{\rho}_{i})+\left.\frac{dU^{X}}{d\rho}\right|_{\bar{\rho}_{i}}(\rho_{i}-\bar{\rho}_{i})+\frac{1}{2}\left.\frac{d^{2}U^{X}}{d\rho^{2}}\right|_{\bar{\rho}_{i}}(\rho_{i}-\bar{\rho}_{i})^{2}\right\rangle_{\mathcal{S}}$
$=U^{X}(\bar{\rho}_{i})+\left.\frac{dU^{X}}{d\rho}\right|_{\bar{\rho}_{i}}\left\langle\rho_{i}-\bar{\rho}_{i}\right\rangle_{\mathcal{S}}+\frac{1}{2}\left.\frac{d^{2}U^{X}}{d\rho^{2}}\right|_{\bar{\rho}_{i}}\left\langle(\rho_{i}-\bar{\rho}_{i})^{2}\right\rangle_{\mathcal{S}}$
$=U^{X}(\bar{\rho}_{i})+\left.\frac{dU^{X}}{d\rho}\right|_{\bar{\rho}_{i}}(\langle\rho_{i}\rangle_{\mathcal{S}}-\bar{\rho}_{i})+\frac{1}{2}\left.\frac{d^{2}U^{X}}{d\rho^{2}}\right|_{\bar{\rho}_{i}}\left(\langle\rho_{i}^{2}\rangle_{\mathcal{S}}-2\langle\rho_{i}\rangle_{\mathcal{S}}\bar{\rho}_{i}+\bar{\rho}_{i}^{2}\right)$
$=U^{X}(\bar{\rho}_{i})+\frac{1}{2}\left.\frac{d^{2}U^{X}}{d\rho^{2}}\right|_{\bar{\rho}_{i}}{\rm Var}_{\mathcal{S}}(\rho_{i}).$ (24)
In passing to the last line, we have used $\langle\rho_{i}\rangle_{\mathcal{S}}=\bar{\rho}_{i}$, and the definition of the variance,
${\rm Var}_{\mathcal{S}}(\rho_{i})=\langle\rho_{i}^{2}\rangle_{\mathcal{S}}-\langle\rho_{i}\rangle_{\mathcal{S}}^{2}.$ (25)
In the event that the second-order term is small compared with the leading term in Eq. (24) (as for example shown for Cu-Ni alloys in [najafabadi:wang:1991]) and ${\rm Var}_{\mathcal{S}}(\rho_{i})$ is small (indicating convergence of the expansion in Eq. (22), then
$\left\langle U^{X}(\rho_{i})\right\rangle\approx U^{X}(\bar{\rho}_{i}),$ (26)
which is the usual assumption in the AA formalism. The average energy is then
$\langle E\rangle=\sum_{i}U_{\rm eff}(\bar{\rho}_{i}),$ (27)
where
$U_{\rm eff}(\bar{\rho})=\sum_{X}U^{X}(\bar{\rho})c_{X},$ (28)
and $\bar{\rho}_{i}$ is given in Eq. (23).
Computationally, this term is exactly as fast to evaluate as the original embedding energy formulation, assuming that the functional forms of $U^{X}(\bar{\rho})$ and $\rho^{Y}_{ij}(r)$ are stored as a discrete grid of points and interpolated on-the-fly at run time. If so, the sum over species $Y$ in Eq. (23) and $X$ in Eq. (28) can be pre-computed for a given set of species concentrations and stored on the same grids of $\bar{\rho}$ and $r$ values.
II.4 Considerations for Dynamic Simulations
To use LOAA in a dynamic setting, we must consider an appropriate way to assign effective masses to the LOAA particles. For this purpose, similar to [dupuy:tadmor:2005], we argue that the LOAA system should have the same total mass and the same momentum free energy as the original, true species system. The momentum free energy, $F_{p}$, of a system of $N$ particles at temperature $T$ is
$F_{p}=-kT\ln\prod_{i=1}^{N}\left(\frac{2kT\pi\mu_{i}}{h_{p}^{2}}\right)^{3/2},$ (29)
where each particle mass $\mu_{i}$ must be that of its species, denoted $m_{X}$ for species $X$, $k$ is the Boltzmann constant, and $h_{p}$ is an arbitrary constant with units of momentum. Assigning masses based on the species, this becomes
$F_{p}=-kT\ln\prod_{X}\left(\frac{2kT\pi m_{X}}{h_{p}^{2}}\right)^{3c_{X}N/2},$ (30)
where $X$ runs over the unique species $A,B,C,\ldots$, and $c_{X}N=N_{X}$ is the number of particles of species $X$. The goal now is to set the mass of all LOAA particles to one value, M, such that we recover this same $F_{p}$ from the LOAA system. The LOAA momentum free energy is
$F^{\rm LOAA}_{p}=-kT\ln\left(\frac{2kT\pi M}{h_{p}^{2}}\right)^{3N/2},$ (31)
where $M$ is the mass of every LOAA particle. Setting $F_{p}=F_{P}^{LOAA}$ requires the mass $M$ to be
$M=\prod_{X}\left(m_{X}\right)^{c_{X}}.$ (32)
This is simply the geometric average of the mass of all the particles in the original system.
III Results
To explore the capabilities of the LOAA methodology, we study two example problems: 1) a 2D binary alloy modeled using an LJ potential where the degree of LO can be tuned by the choice of LJ parameters to allow for systematic comparison with the AA approach; and 2) two 3D metallic alloy systems (FeNiCr and NiAl) modeled using EAM potentials where the effect of LO on mechanical properties and phase transformation is explored.
In what follows, we refer to calculations using standard IPs with the actual atomic species as TS to differentiate with AA and LOAA simulations that employ an effective potential without species identities.
III.1 2D Lennard-Jones Binary Alloy
We begin our study of LOAA with a 2D binary alloy problem. We first consider a random alloy to determine the necessary system size for determining RDFs. Next, we study the effect of LJ parameters on the ordering seen within this alloy. We compare the performance of the standard AA method with the LOAA method to examine the ability of LOAA to capture LO effects. The potential energy per atom and elastic constants of the 2D binary alloy are computed and compared for the TS, AA, and LOAA methods.
III.1.1 2D Lennard-Jones Hexagonal Crystal
Consider a 2D binary alloy of species AA and $B$, with concentrations $c_{A}$ and $c_{B}$. We focus on alloys that are majority AA with small amounts of $B$. The atoms interact via a so-called Kob-Andersen potential that consists of a set of three LJ pair potentials with different parameters for each species combination [kob:andersen:1994, kob:andersen:1995a, kob:andersen:1995b]:
$V_{XY}(r)=4\epsilon_{XY}\left[\left(\frac{\sigma_{XY}}{r}\right)^{12}-\left(\frac{\sigma_{XY}}{r}\right)^{6}\right],$ (33)
where $\sigma_{XY}$ and $\epsilon_{XY}$ are LJ parameters for interaction between species $X$ and $Y$ with units of length and energy, respectively, and $X,Y\in\{A,B\}$. For simplicity, we set all $\sigma$ parameters to be the same, $\sigma_{AA}=\sigma_{BB}=\sigma_{AB}=\sigma$, and the potential is truncated at a cutoff radius $r_{\rm cut}$ for all interactions ($AA$, BBBB, ABAB) and shifted to have zero energy at the cutoff, so we have
$V_{XY}(r)=4\epsilon_{XY}\varphi(r),\qquad\varphi(r)=\hat{\varphi}(r)-\hat{\varphi}(r_{\rm cut}),\qquad\hat{\varphi}(r)=\left[\left(\frac{\sigma}{r}\right)^{12}-\left(\frac{\sigma}{r}\right)^{6}\right].$ (34)
In simulations, we take $\sigma=1$ and $r_{\rm cut}=2.0$.
We study the effect of a range of $\epsilon_{XY}$ values on the LO. In all cases the ground state crystal structure is hexagonal with a nearest-neighbor distance $a$ that depends on the choice of the $\epsilon$ values. For a crystal containing a single species, the result is $a=1.1159\sigma$.
We begin by examining the radial distribution function for the pure A system. Adapting Eq. (48) in Appendix A to a discrete 2D system, the RDF is
$g(r)=\frac{\Delta n(r)}{2\pi r\Delta r(N/\Omega)},$ (35)
where $N/\Omega$ is the number density (number of atoms $N$ per periodic area $\Omega$), $\Delta r=r_{\rm max}/N_{\rm bins}$, $r_{\rm max}$ is the maximum radius computed, $N_{\rm bins}$ is the number of bins in the histogram used in the computation, and $\Delta n(r)$ is the average number of atoms within a distance $r,r+\Delta r]$ of any other atom,
$\Delta n(r)=\frac{1}{N}\sum_{i=1}^{N}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}\mathcal{H}((r_{ij}-r)(r+\Delta r-r_{ij})),$ (36)
where $\mathcal{H}(\cdot)$ is the Heaviside step function. It is important to note that the results of the calculation in Eq. (35) depend on the choice of $\Delta r$ (the smaller $\Delta r$, the larger $g(r)$).
In the results below, $r_{\rm max}=r_{\rm cut}$ and $N_{\rm bins}=100$, so that $\Delta r=2.0/100=0.02$. For a hexagonal crystal, $N/\Omega=(3\times 1/6)/(\sqrt{3}a^{2}/4)=2/(\sqrt{3}a^{2})$, which for $a=1.1159\sigma$ (where $\sigma=1$) is $N/\Omega=0.9273$. The RDF for the pure $A$ crystal is show in Fig. 1. There are three peaks in the range $0,r_{\rm cut}]$ associated with the first, second and third neighbor shells at $r=a=1.1159$, $r=\sqrt{3}a=1.9328$ and $r=2a=2.2318$, which each have 6 neighbors.
Figure 1: RDF for a pure AA hexagonal crystal structure.
Next we consider a binary alloy with concentrations $c_{A}=0.8$ and $c_{B}=0.2$. We begin with the random alloy case where the species of any atom is independent of its environment and is set randomly based on the probabilities defined by the concentration values. Since there are now multiple species, we compute partial RDFs,
$g_{XY}(r)=\frac{\Delta n_{XY}(r)}{2\pi r\Delta r(N_{Y}/\Omega)},$ (37)
where $N_{Y}$ is the number of atoms of species $Y$, and $\Delta n_{XY}(r)$ is the average number of atoms of species $Y$ within a distance $r,r+\Delta r]$ of an atom of species $X$,
$\Delta n_{XY}(r)=\frac{1}{N_{X}}\sum_{i=1}^{N}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}s_{i}^{X}s_{j}^{Y}\mathcal{H}((r_{ij}-r)(r+\Delta r-r_{ij})).$ (38)
Here $s_{i}^{X}$ and $s_{i}^{Y}$ are the species indicator functions defined in Eq. (2).
Figure 2: Normalized partial RDF versus number of atoms in the system at the (a) first, (b) second and (c) third peaks. Calculations performed for system sizes of $4\times 2$ ($N=16$), $16\times 9$ ($N=288$), $32\times 18$ ($N=1152$), $48\times 28$ ($N=2688$), and $64\times 37$ unit cells ($N=4736$), with 2 atoms per cell.
As shown in Appendix A, for the case where the species of the atoms are set independently, we have $g_{XY}(r)=g(r)$. This requirement is helpful in determining a minimal system size to ensure that a single snapshot provides good statistics as assumed in Eq. (14), which leads to the effective potential definition in Eq. (16) that we will use below. Fig. 2 presents $g_{XY}(r)/g(r)$ at the three peaks for different system sizes. We see that the results are largely converged at $N=2688$ atoms, although the $BB$ interactions remain noisier due to the lower concentration of $B$ atoms.
We further perform an energy convergence study to determine the necessary minimum system size as shown in Fig. 3. Increasing from $N=16$ to $N=4736$ atoms, the potential energy per atom is computed for 10 realizations (different random seeds) and standard deviation error bars are plotted in Fig. 2. We again see that the results converge around $N=2688$ atoms. To balance simulation costs with accuracy, the following TS calculations are performed with this system size (2688 atoms).
Figure 3: Potential energy per atom versus system size. Each point is an average over 10 computations with standard deviation error bars. Computed at $(\epsilon_{AB},\epsilon_{BB})=(1.0,0.8)$.
III.1.2 Local Ordering in a 2D LJ Binary Alloy
The propensity to LO depends on the nature of bonding in the material, which in our system is determined by the values of the LJ parameters. The energy of a particular arrangement of $B$ atoms substituted into an $A$ lattice, relative to the energy a pure $A$ lattice, for second-neighbor LJ interactions has the following general form (see Appendix B for the derivation):
$\frac{\Delta E(a_{0};\mathcal{S})}{\epsilon_{AA}}=4C(\mathcal{S})\left[\left(\frac{\sigma}{a_{0}}\right)^{12}-\left(\frac{\sigma}{a_{0}}\right)^{6}\right]+4D(\mathcal{S})\left[\left(\frac{\sigma}{\sqrt{3}a_{0}}\right)^{12}-\left(\frac{\sigma}{\sqrt{3}a_{0}}\right)^{6}\right],$ (39)
where $\mathcal{S}$ is a species arrangement forming an LO pattern (defined by $N_{B}$, the number of $B$ atoms, and the relative positions of the $B$ atoms on the AA lattice), $C(\mathcal{S})$ and $D(\mathcal{S})$ are constants that depend on the number of first- and second-neighbor bonds, respectively, for $AA$, ABAB, and $BB$ interactions, and on the LJ parameter ratios $\epsilon_{AB}/\epsilon_{AA}$ and $\epsilon_{BB}/\epsilon_{AA}$. The LJ length parameters are all set to $\sigma=1$, so that $a_{0}=1.1159$ is the equilibrium lattice constant of a pure $A$ lattice. [^10] Any small relaxations in the vicinity of $B$ atoms that might occur are neglected.
{algorithm}
[t] Calculation of the LO phase diagram for a 2D binary crystal with second-neighbor LJ interactions. (See text for discussion.)
for $\epsilon_{AB}/\epsilon_{AA}\in[0,5]$ do
  for $\epsilon_{BB}/\epsilon_{AA}\in[0,5]$ do
   for $N_{B}=2,\dots,7$ do
     Determine the set $\mathcal{B}(N_{B})$ of unique LO patterns for $N_{B}$
     for all $\mathcal{S}\in\mathcal{B}(N_{B})$ do
      Compute the coefficient $C(\mathcal{S})$ and $D(\mathcal{S})$ using Eq. (65).
      Compute the relative energy $\Delta E(a_{0};\mathcal{S})/\epsilon_{AA}$ using Eq. (39).
     end for
   end for
   Store the minimal relative energy and associated LO patterns $\mathcal{S}$ identified $\forall\mathcal{S}\in\ \mathcal{B}(2)\dots\mathcal{B}(7)$
  end for
end for
Figure 4: Distinct LO patterns for (a) $N_{B}=2$, and (b) $N_{B}=3$.
To explore how the LJ parameters affect LO, we perform a systematic analysis of Eq. (39) identifying the ground state LO patterns across a range of LJ $\epsilon$ parameters. We follow the procedure in Algorithm III.1.2. For each value $\epsilon_{AB}/\epsilon_{AA}$ and $\epsilon_{BB}/\epsilon_{AA}$, LO patterns are sought for $N_{B}=2$ to 7 $B$ atoms. [^11] The LO patterns are constructed on a template of a hexagonal lattice centered on (0,0) with all points up to the third-neighbor distance from the origin, i.e. 18 neighboring sites plus the origin site for a total of 19. The upper limit of $N_{B}=7$ is the maximum number of $B$ atoms that can separate on the template beyond the second-neighbor distance (i.e. separate beyond interaction), which allows all possible configurations to be fully explored. For each value of $N_{B}$, all distinct LO pattern permutations (accounting for hexagonal symmetry) up to third-neighbors are generated. For example, for $N_{B}=2$ there are three distinct LO patterns where the atoms are separated by the first-neighbor distance ($a_{0}$), second-neighbor distance ($\sqrt{3}a_{0}$), and third-neighbor distance ($2a_{0}$); for $N_{B}=3$ there are 9; and so on up to 70 configurations for $N_{B}=7$ (see Fig. 4). For each LO pattern $\mathcal{S}$, the coefficients $C(\mathcal{S})$ and $D(\mathcal{S})$ in Eq. (39) are computed using Eq. (65) based on the number of $AA$, ABAB, and $BB$ bonds in the configurations and specified values of $\epsilon_{AB}/\epsilon_{AA}$ and $\epsilon_{BB}/\epsilon_{AA}$.
Figure 5: Phase diagram and associated simulation results using different $\epsilon$ values. (Left) phase diagram with the green region corresponding to phase separation (PS), the orange region corresponding to a homogeneous solid solution ($S$), and the black line corresponding to a random alloy; (middle) LO patterns; (right) MC simulation results for a 288 atom system with concentrations $c_{A}=0.8$ and $c_{B}=0.2$. Visualization using OVITO [ovito].
Fig. 5 (left) shows the resulting “phase diagram” where each point is colored according to the LO pattern corresponding to the ground state. Interestingly, there are only two distinct LO states in the ground state of this 2D binary system. Each point colored green favors $BB$ interactions leading to phase separation in the ground state in which clumps of $B$ atoms form inside the AA lattice (see Fig. 5 (middle)). In the orange region, $BB$ interactions are unfavorable, and a solid solution is formed with $B$ atoms uniformly distributed in the AA lattice with spacing between them that exceeds their range of interaction (see Fig. 5 (middle)). We will refer to these two regions as the phase separation (PS) and solid solution (SS) regions. Along the black line separating the two regions, the energy of all LO patterns are the same and therefore the $B$ atoms arrange randomly in the ground state — this is the ideal random alloy case. This line where no local-ordering occurs (i.e. the ‘no-LO line’) is defined by the equation:
$\frac{\epsilon_{BB}}{\epsilon_{AA}}=2\frac{\epsilon_{AB}}{\epsilon_{AA}}-1.$ (40)
Substituting this equation into the expressions for $C(\mathcal{S})$ and $D(\mathcal{S})$ in Eq. (65), the result is:
$C(\mathcal{S})$ $=\left(\frac{\epsilon_{AB}}{\epsilon_{AA}}-1\right)\left(N_{AB}^{(1)}+2N_{BB}^{(1)}\right),$ (41a)
$D(\mathcal{S})$ $=\left(\frac{\epsilon_{AB}}{\epsilon_{AA}}-1\right)\left(N_{AB}^{(2)}+2N_{BB}^{(2)}\right).$ (41b)
In addition due to the geometry of the lattice, [^12] In a hexagonal lattice, each $B$ atom has 6 first neighbors and 6 second neighbors, which in total are the sum of AA neighbors plus twice the number of $B$ neighbors since those bonds are shared between $B$ atoms. we have that $N_{AB}^{(1)}+2N_{BB}^{(1)}=N_{AB}^{(2)}+2N_{BB}^{(2)}=6N_{B}$, so that
$C(S)=D(S)=6N_{B}\left(\frac{\epsilon_{AB}}{\epsilon_{AA}}-1\right).$ (42)
This means that at any point on the no-LO line for a given number of $B$ atoms the energy of every LO pattern is the same. Physically this happens because, as seen in Eq. (40), along this line the energy of a $BB$ bond is equal to the energy of two $AB$ bonds, so that neither LO pattern is favorable over the other (the remaining “1-1” term is simply accounting for the relative energy, as described in Appendix B.)
III.1.3 Local Ordering Effect on Ground State RDF and Energy
To test the predictions of the phase diagram in Section III.1.2, we perform Monte Carlo (MC) simulations for the true species (TS) case (i.e. using the the species-dependent IP in Eq. (33)) to determine the ground state configuration for a 2D binary hexagonal crystal supercell. We begin with a crystal of 2688 all $A$ atoms with the equilibrium lattice constant of 1.1159. A specified concentration of $B$ atoms is then distributed onto the lattice. Simulations are performed for three different different choices of $\epsilon$ parameters corresponding to the two LO states and one no-LO state, and $\sigma=1$. To find the ground state configuration, MC swaps of atomic species are performed with a Metropolis acceptance criteria over a series of decreasing temperatures. 320,000 MC swaps are performed at each temperature, starting at $T=4.0$ (LJ temperature units) and halving each time until a final value of $T=1.953125\times 10^{-3}$ (see Supplementary Material (SM)). No relaxation is performed following MC swaps, i.e. the ideal all-$A$ crystal structure is held fixed. Since relaxation effects are small, this procedure leads to a good approximation for the true ground state. All simulations are performed using the LAMMPS package [lammps].
The analysis in Section III.1.2 identifies three distinct regions in $\epsilon$ space that control the resulting ground state crystal structure as seen in Fig. 5 (right). $\epsilon$ values from the PS region of the phase diagram favor nearest-neighbor $B$ bonds, and as may be expected this leads to a ground state where the $B$ atoms clump together. Along the black line, no LO patterns are preferred, and the ground state shows a random pattern. Finally, in the SS region bonding between $B$ is unfavored, and in the ground state the $B$ atoms are spread apart.
Beyond visually validating the LO states, the results from the TS simulations are also examined in terms of their radial distribution functions. Partial RDFs for $AA$, ABAB, and $BB$ interactions from each region in the phase diagram are computed for first-, second-, and third-neighbor distances and shown in Fig. 6. The colors correspond to those in the phase diagram (see figure caption). The blue bars correspond to the total RDF of the system. The partial RDF results are averaged over 10 realizations with error bars shown. As expected from the derivation in Appendix A, the black bars of the randomly distributed system are equal to the total RDF. It can be seen that the three LO states exhibit distinct RDF profiles in accordance with their LO states. Interestingly, the partial RDFs for all points within the same region of the phase diagram are found to be equal, therefore Fig. 6 represents all partial RDFs for the ground state configurations of this system. This observation is explained by the no-LO line relation in Eq. (40). To the left of the no-LO line, $\epsilon_{BB}/\epsilon_{AA}>2\epsilon_{AB}/\epsilon_{AA}-1$, which means that there is a preference for $BB$ bonds since these are lower in energy and hence the same clumping of $B$ atoms is observed at all points in the PS region. In the SS region, the effect is opposite with $AB$ bonds preferred over $BB$ and hence the spreading of $B$ atoms.
Figure 6: Partial radial distribution functions at first, second, and third neighbors for (a) AAAA, (b) ABAB (c) BBBB interactions. Calculations performed for system size of $48\times 28$ unit cells ($N=2688$). Bar chart colors correspond to $\epsilon$ values from different regions of the phase diagram: green=PS, orange=$S$, and black=no-LO. The blue bars correspond to the total (non-partial) RDF. Partial RDFs shown are for $(\epsilon_{AB}/\epsilon_{AA},\epsilon_{BB}/\epsilon_{AA})=(1,3),(3,2),(1,1).$
We can further characterize the ordering seen in each LO state through the Warren-Cowley (WC) order parameters [cowley:1965, norman:warren:1951], defined by
$\alpha_{ij}^{XY}=1-\frac{p(s_{i}^{X}=1,s_{j}^{Y}=1\,|\,\mathcal{C})}{c_{X}c_{Y}}$ (43)
for atoms $i$ and $j$ of species $X$ and $Y$, where the probability is defined in Eq. (14). [^13] The standard Warren-Cowley parameter formula involves a conditional probability, but it can trivially be shown that the given formulation with a joint probability is also correct [rao:curtin:2022]. Fig. 7 displays the WC parameters for each atom type combination at nearest neighbor distance. The three LO states are plotted in accordance with their coloring in Fig. 5. A WC parameter that is <0<0 indicates an attractive interaction between atom types, while >0>0 indicates repulsion. A WC parameter equal to 0 suggests that there is no LO, which explains why the no-LO state is not visible in Fig. 7. Only first neighbor WC parameters were plotted as the second neighbor parameters are nearly identical.
Figure 7: Warren-Cowley order parameters for nearest neighbor interactions. Calculations performed for system size of $48\times 28$ unit cells ($N=2688$). Colors correspond to those from phase diagram regions. $(\epsilon_{AB}/\epsilon_{AA},\epsilon_{BB}/\epsilon_{AA})=(1,3),(3,2),(1,1).$
Figure 8: TS energy contours over $\epsilon$-space. The no-LO line is shown in red.
The potential energy per atom across $\epsilon$ space is visualized in the energy contour plot in Fig. 8. The energy is computed at 100 points across phase diagram space. At each point, results are averaged over 10 ground state realizations obtained via MC with different random seeds. This plot can be understood by considering the different LO states on the phase diagram. In the SS region of Fig. 5, $B$ atoms are not (or are minimally) interacting, which explains the vertical energy contour lines in this region since there is little dependence on $\epsilon_{BB}$. In the PS region of the phase diagram $B$ atoms form clumps, and here we see that the energy has little dependence on $\epsilon_{AB}$. This makes sense since there are fewer $AB$ interactions as these are only present along the borders of the clumps of $B$ atoms.
III.1.4 AA and LOAA Predictions for Ground State Energy
As seen from the TS analysis, multiple realizations of large and computationally expensive MC simulations are necessary to obtain reliable statistical estimates for the ground state properties. As explained in Section II, the AA and LOAA methods directly predict statistical averages with smaller supercells and without the need for repeated simulations. The AA approach assumes a random alloy neglecting the effect of LO, and LOAA includes LO effects via the partial RDFs of the ground state.
The AA potential form is given in Eqns. (8) and (9). For the the special case of a LJ pair potential where the $\sigma$ parameter for all interactions are the same, the effective AA potential energy has the simple form
$V_{\rm eff}(r)=4\bar{\epsilon}\varphi(r),$ (44)
where $\varphi(r)$ is defined in Eq. (34) and
$\bar{\epsilon}=\epsilon_{AA}c_{A}^{2}+2\epsilon_{AB}c_{A}c_{B}+\epsilon_{BB}c_{B}^{2}.$ (45)
This can be used with the standard LJ implementation in LAMMPS.
The LOAA potential is defined in Eqns. (16) and (17). The partial RDFs used in Eq. (17) are averaged over 10 different random seeds at 100 individual points across $\epsilon$ space using the TS simulations described in Section III.1.3. To implement the LOAA formulation, we create a new pair stlye in LAMMPS. The LOAA potential is created by modifying the existing LJ pair style to include the additional $G_{XY}(R)$ term seen in Eq. (16).
Figure 9: Difference in the energy per atom between the (a) AA method, and (b) LOAA method, and the exact TS results plotted across $\epsilon$-space. The no-LO line is shown in red.
The performance of the AA and LOAA is evaluated by comparing their predictions for the potential energy per atom with the exact results obtained using TS simulations across $\epsilon$-space.
The results for the AA method are presented in Fig. 9(a). As expected, we see that the method is exact only in the case of no LO (shown as a red line in the plot). The error in the AA energy increases linearly with the distance from this line as shown by a theoretical analysis in Appendix C. Comparing with the TS energy contour plot in Fig. 8, we find that the relative error in the AA energy prediction can be as large as 45% depending on the location in $\epsilon$-space (see Fig. S2 in the SM). Since, as explained in Section III.1.3, the partial RDFs are the same at all points in a given region of the phase diagram, the increasing error in AA method predictions is due to an increase in the $BB$ bond strength rather than an increase in the degree of LO.
We note that errors are larger to the left of the no-LO line (corresponding to the PS region in Fig. 5) relative to the right (SS region). Referring to the energy difference expression in Eq. (76), we see that the error in the direction $t$ perpendicular to the no-LO line scales with with $\zeta-2\gamma$. Due to the nature of the partial RDF profiles in the PS and SS regions, this factor is considerably larger in magnitude (around four times) in the PS region than in the SS region. A systematic increase in error occurs on both side of the no-LO line, but this explains why the error grows more rapidly with distance to the left of the no-LO line.
The results for the LOAA method are presented in Fig. 9(b) using the same contour scale as for the AA method in Fig. 9(a). The maximum error in energy for LOAA is about $6\times 10^{-3}$, not discernible on this scale. This small error reflects finite size limitations of the TS simulation rather than the LOAA formalism, which is exact. Both the partial RDFs and the energy per atom computed from the TS simulations were obtained from 10 realizations of a 2688 atom system, and as seen in Figs. 2 and 3, neither are fully converged. As a result both the TS energy and the LOAA-predicted energy (which depends on the partial RDFs) have errors and do not match exactly.
III.1.5 AA and LOAA Predictions for Elastic Constants
To further validate the LOAA method, elastic constants are computed at various points in $\epsilon$-space using the TS, AA, and LOAA approaches. Similar to the ground state energy studies, we expect the AA method to be correct only for points on the no-LO line. In contrast, the LOAA method should yield accurate elastic constants for all $\epsilon$ values as it accounts for LO.
The elasticity tensor $\bm{c}$ is computed using the expression for pair potentials from [tadmor:miller:2011, Section 11.5.2]:
$c_{ijkl}=\frac{1}{2\Omega}\sum_{\begin{subarray}{c}\alpha,\beta\\ \alpha\neq\beta\end{subarray}}\left[V_{\text{eff},\alpha\beta}^{\prime\prime}(r^{\alpha\mathring{\beta}})-\frac{V_{\text{eff},\alpha\beta}^{\prime}(r^{\alpha\mathring{\beta}})}{r^{\alpha\mathring{\beta}}}\right]\frac{r_{i}^{\alpha\mathring{\beta}}r_{j}^{\alpha\mathring{\beta}}r_{k}^{\alpha\mathring{\beta}}r_{l}^{\alpha\mathring{\beta}}}{(r^{\alpha\mathring{\beta}})^{2}},$ (46)
where $\Omega$ is the simulation cell volume, and the sum is over all atoms $\alpha$ and $\beta$ in the simulation cell, $\mathring{\beta}$ refers to the nearest periodic copy of atom $\beta$ assuming the minimum image convention, and $r^{\alpha\mathring{\beta}}$ is the distance between atoms $\alpha$ and $\mathring{\beta}$. Because the system is isotropic and 2D, the elasticity tensor has two independent elastic constants, which in Voigt notation are denoted $C_{11}$ and $C_{12}$.
Figure 10: (a) $\text{C}_{11}$ and (b) $\text{C}_{12}$ as a function of perpendicular distance from the no-LO line, $t$, computed using the TS, AA, and LOAA methods. Standard deviation error bars are plotted for TS.
Given the observation in Section III.1.4 that the error in AA prediction scales with the perpendicular distance $t$ from the no-LO line, we perform a similar analysis for the elastic constants. The results using each of the three methods (AA, LOAA, TS) are plotted in Figs. 10(a) and 10(b). The plots are nearly identical at all points along the no-LO line (i.e. at all values of $s$). We see that LOAA predictions agree precisely with the TS ground truth, whereas the AA predictions diverge in proportion to the distance from the no-LO line. The effect is particularly noticeable to the right of the no-LO line ($t>0$) corresponding to the SS region where there is a change in slope that AA fails to capture. This change in behavior is explained by an analytical derivation of the AA and LOAA elastic constants expressions in Appendix D. This analysis provides the scaling factors of the AA and LOAA elastic constants with distance from the no-LO line (i.e. the slopes in Figs. 10(a) and 10(b)). For example for $C_{11}$ the slopes are
$m^{\rm AA}_{11}$ $=\frac{2\epsilon_{AA}}{\sqrt{5}\Omega}\left(c_{B}^{2}-4c_{A}c_{B}\right)\sum_{\begin{subarray}{c}\alpha,\beta\\ \alpha\neq\beta\end{subarray}}\left[\varphi^{\prime\prime}(r^{\alpha\mathring{\beta}})-\frac{\varphi^{\prime}(r^{\alpha\mathring{\beta}})}{r^{\alpha\mathring{\beta}}}\right]\frac{\left(r_{1}^{\alpha\mathring{\beta}}\right)^{4}}{(r^{\alpha\mathring{\beta}})^{2}},$ (47a)
$m^{\rm LOAA}_{11}$ $=\frac{2\epsilon_{AA}}{\sqrt{5}\Omega}\sum_{\begin{subarray}{c}\alpha,\beta\\ \alpha\neq\beta\end{subarray}}\left(c_{B}^{2}\Phi_{BB}(r^{\alpha\mathring{\beta}})-4c_{A}c_{B}\Phi_{AB}(r^{\alpha\mathring{\beta}})\right)\left[\varphi^{\prime\prime}(r^{\alpha\mathring{\beta}})-\frac{\varphi^{\prime}(r^{\alpha\mathring{\beta}})}{r^{\alpha\mathring{\beta}}}\right]\frac{\left(r_{1}^{\alpha\mathring{\beta}}\right)^{4}}{(r^{\alpha\mathring{\beta}})^{2}}.$ (47b)
We see that the slope for the AA method depends only on the concentrations $c_{A}$ and $c_{B}$ and the crystal structure and so will be the same in Regions I and II as seen in Fig. 10(a). In contrast the LOAA slope depends on the partial RDFs through the probability functions $\Phi_{AB}(r)$ and $\Phi_{BB}(r)$ (see Eq. (15)), which explains the change in slope across the no-LO line.
III.2 3D EAM Alloys: FeNiCr and NiAl
As further tests and demonstrations of the LOAA method, we also explore two 3D EAM alloys. First, we consider the ordering effects on lattice constant, cohesive energy, and elastic properties in an FeNiCr alloy as described by the potentials of Bonny et al. [Bonny_2011]. This is the same system studied by [varvenne:luque:2016] for random alloys. We then consider temperature and stress dependent martensitic phase transformations in NiAl using the potential of [pun:mishin:2009].
All 3D simulations are performed with LAMMPS and visualized using Ovito.
III.2.1 Ordering Effects in FeNiCr
In order to directly compare to the AA results of [varvenne:luque:2016], we consider alloys of the composition
Fe(1-x)/2Ni(1-x)/2Crx for various Cr concentrations $x$. As in the previous sections, we will denote the results of the full EAM description with atoms assigned their true species as “TS”. These can either be “random”, in which case atoms are randomly assigned a species consistent with the concentrations, or “ordered”, in which case the species of each lattice site is determined by MC as described below. All TS results are shown for a system of 32,00032{,}000 atoms on $20\times 20\times 20$ units cells of an fcc crystal with periodic boundary conditions.
The ordered TS sample for each unique set of concentrations is determined by an on-lattice MC anneal of the 32,00032{,}000 atoms, initially assigned random species on an fcc lattice with the relaxed lattice constant determined by a separate minimization of the randomized species. The MC annealing then proceeds to attempt 100,000100{,}000 species swaps at each of 6 progressively lower numerical temperatures ($1200K$, $1000K$, $800K$, $600K$, $400K$, $200K$).
Figure 11: Comparison of TS, AA and LOAA calculations for Fe(1-x)/2Ni(1-x)/2Crx alloys. (a) Lattice constant, (b) cohesive energy, (c) and (d) elastic properties.
Results of the AA and LOAA formulations are compared to the TS results in Fig. 11. Note that both the AA and LOAA simulations can be performed on minimal periodic simulation cells (these were done on 32-atom fcc crystals). Since the formulations homogenize the atoms to make them all equivalent, there is no longer a need to represent the complex ordered (or disordered) state or to have enough atoms to accurately match the concentrations of each species. The additional requirement of the LOAA formulation — the function $G(R)$ — is determined from the RDFs of the final configuration of the same on-lattice MC anneal described above.
Figure 12: Comparison of TS, AA and LOAA elastic constants when the TS results are not allowed to relax from perfect lattice positions.
First, we compare the two TS lines (random and ordered) in each frame of Fig. 11. The difference between these curves is the effect of LO, which is not insignificant in these alloys. The AA formulation does an excellent job of capturing the lattice constant and cohesive energy for the random TS, and a reasonable job of the random TS elastic constants. AA is of course unable to capture the ordered TS curves, since it knows nothing of the LO. On the other hand, we see that the LOAA results are in excellent agreement for $a_{0}$ and $E_{\rm coh}$ case, and has comparable error to that of AA when considering the elastic constants. The elastic constants are only in modest agreement with the TS results (for both AA and LOAA). This is due to relaxation effects in the TS results, which arise from local variations of atomic environments. Relaxation effects are absent from the homogenized AA and LOAA results. In Fig. 12, we compare AA and LOAA to TS results where the atoms are not allowed to relax from their perfect crystal positions, and we see the agreement largely restored.
III.2.2 Radial Distribution Functions for NiAl
Figure 13: (a)-(c) Partial RDFs for the ordered TS species, compared with the expected RDF for a random distribution of the Ni and Al atoms. (d) resulting $G(R_{0})$ for the Ni0.67Al0.33 alloy.
For the NiAl system, we begin by examining the partial RDFs, the resulting $G(R)$, and the effect these have on the effective pair potential, then in Section III.2.3 we examine NiAl phase transformations.
In Fig. 13, we compare the partial RDFs for the Ni–Ni, Ni–Al and Al–Al pairs in the ordered TS sample after the MC anneal (recall that this configuration is used to generate $G(R)$ for the LOAA effective pair potentials). In each case, we compare to the bcc RDF for the case where the atomic species are randomly assigned, which is just the RDF for ideal bcc ignoring species. These plots emphasize the effect of the ordering. For example, we see virtually no Al–Al bonds at the nearest neighbor distance, compared to the number of Ni–Ni and Ni–Al bonds. The LOAA formulation takes this into account through the $G(R)$ functions, as shown in Fig. 13(d). When considering a bond that is at some distance $R_{0}$ in the reference configuration, its contribution is a weighted sum of the three types of interactions, not only by average overall concentration (as in AA), but also by the relative probability of finding those bonding pairs in the ordered structure.
Figure 14: (a) Comparison of the TS pair potentials with the effective pair potential for the AA formulation, for the Ni0.67Al0.33 case. (b) Further comparison to the LOAA effective pair potentials, one for each of the first 5 non-zero values in the RDF functions.
The effective pair potentials for AA and LOAA are presented in Fig. 14. In AA, $V_{\rm eff}(r)$ is simply a blend of the three inter-species pair potentials, weighted by their respective compositions. In LOAA, each peak in the RDF curves, corresponding to a different reference configuration pair distance, $R_{0}$, produces a different effective pair potential, through the effect of $G(R)$. This is seen in Fig. 14(b), where the LOAA potential for the first five neighbor shell distances in the reference configuration are compared to the original EAM potentials and the blended result from AA. This highlights the pronounced differences between the AA and LOAA formulation, allowing LOAA to predict the stabilization of the B2 phase at high temperature.
III.2.3 Phase Transformations in NiAl
Perhaps a more interesting and rigorous test of LOAA is the case of martensitic phase transformations in NiAl. Using the EAM potentials they developed [pun:mishin:2009], Pun and Mishin showed [pun:mishin:2010] that the B2 lattice is stable at high temperatures for a variety of off-stochiometric NiAl compositions, provided that the atomic species are determined through an appropriate MC annealing process (i.e. they are not simply random). Further, many of these compositions undergo a martensitic phase transformation to a face-centered tetragonal (fct) structure upon cooling for certain applied stresses. In the following we focus on the specific case of Ni0.67Al0.33   and consider, somewhat arbitrarily simply as an example, its phase transformations upon cooling from 900K to 100K and subsequent re-heating from 100K to 900K while subjecting it to a compressive stress of 900MPa along the $[001]$ lattice direction of the B2 structure.
Such phase transformations will be, to some extent, stochastic in nature, and are therefore sensitive to the initial structure, initial velocities, system size, heating/cooling rate, and temperature control algorithm. Here, we make convenient choices and keep them consistent across all samples (with some clear exceptions explained below). Thus, we are comparing models (TS vs AA and LOAA) as opposed to trying to accurately predict transformation temperatures. Unless otherwise specified, we use a block of 6000 atoms in the B2 phase with an initial lattice constant of 2.8 Å. The box is initially orthogonal with the $x$, $y$ and $z$ axes oriented along $\bar{1}0] [1¯1$0][$\bar${1}10] and $[001]$, respectively. This is run using a Nosé–Hoover controlled NPT ensemble equilibrated at 900 K and with all stress components set to zero except the applied stress for 10,000 timesteps ($\Delta t=1$ fs). After this, the temperature is linearly ramped down from 900K to 100K over 500,000 steps, and then ramped back up to 900 over another 500,000 steps. This simulation allows the simulation box to change size and shape in order to maintain the correct stresses (it is a triclinic box in the language of LAMMPS, defined by three non-orthogonal periodic vectors). In the images of the simulation cell that follow, we show the $[001]$ or $z$ direction as “up” unless the image is specified as a TOP view, in which case the view is directly along the $z$-direction.
By monitoring the length of the simulation box along the direction of the applied stress versus the temperature, we can capture the hysteresis loop of the phase transformations. Upon cooling, the length will contract slowly at first, due to simple thermal effects. It will then contract suddenly at a critical temperature called the martensite start temperature, $M_{s}$. Upon heating, there is again a sudden change ending on the original cooling line at the austenite finish temperature, $A_{f}$. Because $A_{f}$ is higher than $M_{s}$, a hysteresis loop is seen. This can be seen by skipping ahead to Fig. 16.
Figure 15: (a) Initial structure of the random TS Ni0.67Al0.33 sample, colored by species (red is Ni, blue is Al). (b) The same structure as (a) colored by Ovito’s CNA, with blue indicating the B2 structure. (c) The final structure after a spontaneous phase transformation because the B2 structure is unstable. Here green is fcc-like and red is hcp-like, showing twin planes in the fcc-like fct crystal. (d) TOP view of (c) to show the degree of non-orthogonality after the transformation.
First, we briefly consider the TS case where Ni and Al species are randomly assigned to a B2 lattice, as shown in Figs. 15(a) and (b). In (b), the atoms are coloured according to OVITO’s Common Neighbor Analysis (CNA) tool, which detects nominally bcc, fcc, and hcp local atomic arrangements. The B2 stucture of Ni0.67Al0.33 appears as nominally bcc (blue) using this tool. In this case, the B2 structure is not stable, and the crystal spontaneously transforms to a twinned fcc-like structure, seen in green and red in Fig. 15(c) and (d), where the twin planes appear as locally hcp-like and are colored red. The simulation cell remains nearly orthogonal during the transformation, but contracts along $z$ and expands along $x$ and $y$. This illustrates the importance of the LO (short-range order) in stabilizing the B2 structure. The random placement of the Al within the Ni lattice simply produces a random solid solution for which the bcc arrangement is not stable at any temperature, and the stable structure is fcc, distorted by the disparate sizes of the Ni and Al atoms.
Figure 16: Cooling (blue) and heating (red) for the ordered TS Ni0.67Al0.33 sample, showing the temperatures $M_{s}$, $A_{f}$ and the hysteresis loop.
Figure 17: (a) Initial structure of the ordered TS Ni0.67Al0.33 sample, colored by species (red is Ni, blue is Al). (b) The same structure as (a) colored by Ovito’s CNA, with blue indicating the B2 structure. (c) The final structure after the cooling phase, which has transformed to a twinned fct structure. (c) The final structure upon reheating, restored to the original B2 with a few thermal fluctations.
Next, the TS case is run again, but this time with the atomic species set by an MC annealing process identical to the one described above for the FeNiCr system. The arrangement of the atomic species stabilizes the B2 structure at high temperature. The resulting hysteresis loop and structures are shown in Figs. 16 and 17. We take these plots to be the “ground truth” for the Ni0.67Al0.33 system — it is the curve we hope to reproduce with our averaged-atom models. In Fig. 17, we see that the crystal indeed transforms from the initial B2 in (a) and (b), to fct upon cooling in (c), and back to B2 after re-heating in (d). The fact that the low temperature phase is fct, and not fcc, is not obvious from the images, but can be verified by analyzing the lattice vectors carefully.
We now run the same simulation, but with the average-atom interatomic model (both AA and LOAA). We start with the same 6000 atom simulation box described above, but now all atoms are equivalent averaged-atoms instead of Ni or Al. We set their masses to be the geometric average based on the concentrations, i.e. $M=m_{Ni}^{0.67}m_{Al}^{0.33}=45.42$ amu, as per Section II.4.
Figure 18: (a) Final structure of the AA simulation after a spontaneous phase transformation because the B2 structure is unstable (compare to Fig. 15(c). (b) Alternate TOP view to highlight non-orthogonal box. (The initial structure was the same as Fig. 15(b).)
Figure 19: Cooling (blue) and heating (red) for the LOAA Ni0.67Al0.33 sample, showing the temperatures $M_{s}$, $A_{f}$ and the hysteresis loop.
Figure 20: All images are TOP view, looking along $z$. (a) Transformed structure of the LOAA simulation at the end of the cooling phase (the initial high temperature phase was the same as Fig. 17(b)). In (b), we see that the structure is still fct martensite at about 620K during heating, just before the first rise in the heating curve. (c) The structure at about 700K, just after the first rise, which we see by the direction of the shear is still fct, but a different variant. (d) The final structure at 900K, which has transformed back to B2.
Using the AA formulation produces essentially the same behavior as the random TS sample, i.e. there is a spontaneous transformation to the fcc phase because the B2 phase is unstable for the random alloy. This is shown in Fig. 18. On the other hand, the LOAA formulation shows the same B2 stability at high temperatures and the phase transformation as the ordered TS sample, albeit with different transformation temperatures. This can be seen in Figs. 19 and 20. The hysteresis loop for LOAA (Fig. 19) shows an unusual double rise on heating. The first rise at around 650 K is actually not the phase transition, but rather a jump between two variants of the martensite before the final true transformation to the austenite at around 755 K. This is shown in the snapshots of Fig. 20, where (a) is the transformed martensite after cooling, (b) is the same martensite but at a temperature of around 620K, just before the first abrupt rise in the heating curve. In (c), we see that the crystal is still in the martensitic phase at about 640 K (after the rise), although sheared in the other direction. Finally, (d) shows the crystal after it has fully transformed back to B2 austenite, made clear by the orthogonal simulation box and the blue coloring (the noisy mix of CNA values in (b)–(d) is due to atomic vibrations at high temperature).
IV Summary
An extension to the AA approach accounting for LO (short-range order) is derived based on partial RDF information obtained on a reference configuration. The result is an effective IP defined for an atomistic system without species, which in the limit of no structural relaxation provides exact results for an ensemble over all possible species arrangements on the reference configuration. Once the effective IP is constructed, it has a similar cost to the original “true species” IP upon which it is based, but enables simulations with far smaller supercells and without the need to run multiple times for good statistics for significant computational savings.
LOAA is validated on several example problems. First, a 2D binary hexagonal crystal modeled via LJ interactions. A phase diagram is constructed showing the dependence of the nature and intensity of LO as a function of LJ parameters. It is shown that LOAA accurately reproduces the energy and elastic constants of the true species IP for LO patterns, in contrast to AA that is only correct for the random alloy case where no LO exists. Second, LOAA is shown to accurately reproduce the lattice constant, cohesive energy, and elastic properties of an ordered Fe(1-x)/2Ni(1-x)/2Crx alloy across for $x$ in the range 0 to 0.5 modeled via an EAM potential. Finally, third, a temperature driven martensitic phase transformation under axial compression in an off-stoichiometry Ni0.67Al0.33 alloy modeled via EAM is studied. It is shown that LO is necessary to stabilize the high-temperature B2 phase. Upon cooling the crystal exhibits hysteresis as it transforms to a low temperature fct phase, and then back to B2 upon heating. The LOAA effective IP is able to capture this effect, albeit at somewhat different transformation temperatures, whereas the AA effective destabilizes into a twinned fcc-like structure.
The method can be readily extended to other IP functional forms, and would be of interest to apply to complex materials, like HEAs that require very large supercells to capture the off-stoichiometric structure, and where LO is known to play an important role. On the methodological side, the approach can be extended to account for local relaxation effects by considering the effect of such relaxation on the partial RDF form. Preliminary work along these lines appears promising.
Code and example input suitable for running LOAA in LAMMPS is provided in the SM.
Acknowledgements.
CZ and ET acknowledge support by NASA Space Technology Graduate Research Opportunity (NSTGRO) Award 21-NSTGRO22-0159. RM acknowledges the support of the Natural Sciences and Engineering Research Council of Canada (NSERC). The authors thank Ryan S. Elliott for helpful discussions, and Lauren Abbott and the computational materials group at NASA Ames Resarch Center for their valuable assistance.
Appendix A Radial Distribution Functions
The radial distribution function for a monoatomic system (or when not distinguishing atom species) is
$g(r)=\frac{dn(r)}{4\pi r^{2}dr(N/V)},$ (48)
where $N/V$ is the number density (number of atoms $N$ per volume $V$), and $dn(r)$ is the average number of atoms within a distance $[r,r+dr]$ of any other atom,
$dn(r)=\frac{1}{N}\sum_{i=1}^{N}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}\langle\delta_{\rm D}(r_{ij}-r)\rangle,$ (49)
where $\delta_{\rm D}()$ is the Dirac delta function, and $\langle\cdot\rangle$ indicates the canonical phase average at a given temperature $T$.
Formally, the RDF in the canonical ensemble is independent of structure. This is because phase averages are over all configurations (subject to an overall volume constraint), which corresponds to a liquid state. The dependence on structure comes in once the ensemble is restricted in the sense described in Tadmor and Miller [tadmor:miller:2011] to a subset of configuration space, e.g. a face-centered cubic (fcc) arrangement with perturbations about mean positions. In this case, the RDFs will depend on the restriction. In practice, $g(r)$ computed in a molecular dynamics simulation or measured in an experiment is inherently restricted by the accessible time scale.
In the multiple species case, the partial RDF is
$g_{XY}(r)=\frac{dn_{XY}(r)}{4\pi r^{2}dr(N_{Y}/V)},$ (50)
where $N_{Y}$ is the number of atoms of species $Y$, and $dn_{XY}(r)$ is the average number of atoms of species $Y$ within a distance $[r,r+dr]$ of an atom of species $X$,
$dn_{XY}(r)=\frac{1}{N_{X}}\sum_{i=1}^{N}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{N}s_{i}^{X}s_{j}^{Y}\langle\delta_{\rm D}(r_{ij}-r)\rangle.$ (51)
Here $s_{i}^{X}$ and $s_{i}^{Y}$ are the species indicator functions defined in Eq. (2).
The probability of atoms $i$ and $j$ separated by distance $r_{ij}$ being of species $X$ and $Y$, respectively, can be expressed in terms of the numbers $dn_{XY}(r)$ as follows:
$p(s_{i}^{X}=1,s_{j}^{Y}=1\,|\,T)=\frac{c_{X}dn_{XY}(r)}{\sum_{A}\sum_{B}c_{A}dn_{AB}(r)}.$ (52)
The multiplication by the concentrations is required so that for all terms the number of neighbors is normalized by the total number of atoms, and not the number of atoms of a specific species, e.g. for the $dn_{XY}$ term, $c_{X}/N_{X}=c_{X}/(c_{X}N)=1/N$. Next using Eq. (50) to express $dn_{XY}$ in terms of $g_{XY}$ and using $N_{Y}=c_{Y}N$, we have
$p(s_{i}^{X}=1,s_{j}^{Y}=1\,|\,T)=\frac{c_{X}c_{Y}g_{XY}(r)}{\sum_{A}\sum_{B}c_{A}c_{B}g_{AB}(r)}.$ (53)
This is the expression in Eq. (14).
For the special case where the species of the atoms are set independently, we have
$dn_{XY}(r)=c_{Y}dn(r),$ (54)
since the number of neighbors of type $Y$ follows from the concentration $c_{Y}$. Substituting Eq. (54) into Eq. (50) gives
$g_{XY}(r)=\frac{c_{Y}dn(r)}{4\pi r^{2}dr(N_{Y}/V)}=g(r),$ (55)
where we have used $N_{Y}=c_{Y}N$ to obtain the final equality. Thus when the atom species are independent, all partial RDFs are equal to the RDF, which does not account for species. In this case, the probability in Eq. (53) is as expected
$p(s_{i}^{X}=1,s_{j}^{Y}=1\,|\,T)=\frac{c_{X}c_{Y}g(r)}{\sum_{A}\sum_{B}c_{A}c_{B}g(r)}=\frac{c_{X}c_{Y}}{\sum_{A}\sum_{B}c_{A}c_{B}}=c_{X}c_{Y},$ (56)
where we have used $\sum_{A}\sum_{B}c_{A}c_{B}=\sum_{A}c_{A}(\sum_{B}c_{B})=\sum_{A}c_{A}(1)=(1)(1)=1$.
Appendix B Relative bonding energies in 2D Hexagonal Binary Crystals
We derive an expression for the energy of a 2D binary hexagonal crystal with AA and $B$ atoms relative to that of a crystal with all $A$ atoms. Assuming pairwise interactions, the energy of a system of $N$ atoms is given by
$E=\frac{1}{2}\sum_{i}\sum_{j\neq i}\sum_{X}\sum_{Y}V_{XY}(r_{ij})s_{i}^{X}s_{j}^{Y},$ (57)
where $X$ and $Y$ are species, and we take an LJ form with different $\epsilon$ parameter for each pair of species, but the same $\sigma$, so that
$V_{XY}(r)=\epsilon_{XY}\hat{\varphi}(r),$ (58)
where $\hat{\varphi}(r)$ is the LJ function given in Eq. (34). [^14] Note that since no relaxation takes place there is no difference in using $\hat{\varphi}(r)$ rather than the truncated potential in Eq. (34). Normalizing the energy by $\epsilon_{AA}$, the dimensionless energy for a periodic system of $N$ type $A$ atoms on a 2D hexagonal lattice with second neighbor interactions is
$\frac{E(a;\mathcal{S}_{A})}{\epsilon_{AA}}=\frac{1}{2}(6N\varphi(a)+6N\varphi(\sqrt{3}a)),$ (59)
where $a$ is the lattice constant and $\mathcal{S}_{A}$ refers to a species arrangement of all $A$ atoms. For the LJ functional form, this is
$\frac{E(a;\mathcal{S}_{A})}{\epsilon_{AA}}=12N\left[\left(\frac{\sigma}{a}\right)^{12}-\left(\frac{\sigma}{a}\right)^{6}+\left(\frac{\sigma}{\sqrt{3}a}\right)^{12}-\left(\frac{\sigma}{\sqrt{3}a}\right)^{6}\right].$ (60)
Setting $\frac{dE}{da}(a_{0};S_{A})=0$, we find that the equilibrium lattice constant (for second neighbor interactions) is $a_{0}=1.1159$.
Figure 21: 2D hexagonal crystal of AA atoms with two $B$ atoms at a second neighbor distance.
Next, we consider the case of a crystal of AA atoms with two $B$ atoms interacting at a second neighbor distance (denoted $\mathcal{S}_{AB}$) as shown in Fig. 21. In order to obtain a closed-form expression, we neglect the effect of local relaxations about $B$ atoms (which are expected to be small). The energy for this case is derived by considering the number of $AA$, ABAB and $BB$ bonds at first- and second-neighbor distances. The result is
$\frac{E(a;\mathcal{S}_{AB})}{\epsilon_{AA}}=\left(12\frac{\epsilon_{AB}}{\epsilon_{AA}}+3N-12\right)\varphi(a)+\left(10\frac{\epsilon_{AB}}{\epsilon_{AA}}+\frac{\epsilon_{BB}}{\epsilon_{AA}}+3N-11\right)\varphi(\sqrt{3}a).$ (61)
In order to obtain an expression for an infinite system independent of $N$, we consider the energy relative to $E(a;\mathcal{S}_{A})$, i.e. the difference between the energy of the system with $B$ atoms and the energy of the system with all $A$ atoms:
$\frac{\Delta E(a;\mathcal{S}_{AB})}{\epsilon_{AA}}\equiv\frac{E(a;\mathcal{S}_{AB})-E(a;\mathcal{S}_{A})}{\epsilon_{AA}}=\left(12\frac{\epsilon_{AB}}{\epsilon_{AA}}-12\right)\varphi(a)+\left(10\frac{\epsilon_{AB}}{\epsilon_{AA}}+\frac{\epsilon_{BB}}{\epsilon_{AA}}-11\right)\varphi(\sqrt{3}a).$ (62)
The general form of the relative energy for any species arrangement $\mathcal{S}$ containing any number of $B$ atoms at any positions takes the form,
$\frac{\Delta E(a;\mathcal{S})}{\epsilon_{AA}}=C(\mathcal{S})\varphi(a)+D(\mathcal{S})\varphi(\sqrt{3}a).$ (63)
where $C(\mathcal{S})$ and $D(\mathcal{S})$ are dependent on the number of first- and second-neighbor bonds, respectively, for $AA$, ABAB, and $BB$ interactions in a particular species arrangement $\mathcal{S}$.
Finally, for the LJ form, we obtain
$\frac{\Delta E(a;\mathcal{S})}{\epsilon_{AA}}=4C(\mathcal{S})\left[\left(\frac{\sigma}{a}\right)^{12}-\left(\frac{\sigma}{a}\right)^{6}\right]+4D(\mathcal{S})\left[\left(\frac{\sigma}{\sqrt{3}a}\right)^{12}-\left(\frac{\sigma}{\sqrt{3}a}\right)^{6}\right].$ (64)
where $C(S)$ and $D(S)$ are defined by
$C(S)$ $=N^{(1)}_{AB}\frac{\epsilon_{AB}}{\epsilon_{AA}}+N^{(1)}_{BB}\frac{\epsilon_{BB}}{\epsilon_{AA}}-(N^{(1)}_{AB}+N^{(1)}_{BB})$ (65a)
$D(S)$ $=N^{(2)}_{AB}\frac{\epsilon_{AB}}{\epsilon_{AA}}+N^{(2)}_{BB}\frac{\epsilon_{BB}}{\epsilon_{AA}}-(N^{(2)}_{AB}+N^{(2)}_{BB})$ (65b)
where $N^{(1)}_{XY}$ and $N^{(2)}_{XY}$ are the number of $XY$ bonds at nearest neighbor and second neighbor distances, respectively.
By substituting $a=a_{0}$, we obtain a general expression for the minimum energy of a 2D hexagonal crystal with arbitrary arrangement of $B$ atoms.
Appendix C Theoretical Analysis of the Error in Energy Predictions of the AA Method
Fig. 9(a) is a plot of the difference between the energy-per-atom predicted by the AA method and the exact TS case. The plot shows that the energy difference increases as a function of the vertical distance from the no-LO line (shown in red). In this section we derive an analytical result for the energy difference in order to explain this observation. Since an analytical expression for the TS case does not exist (as it involves numerical MC simulations), we use the LOAA solution as a surrogate since it is in close agreement with the TS results.
The effective potentials for AA and LOAA are given in Eq. (9) and (16), respectively. The difference between them $\Delta V_{\rm eff}=V_{\rm AA}-V_{\rm LOAA}$ is
$\Delta V_{\rm eff}(r,R)$ $=\sum_{X}\sum_{Y}V_{XY}(r)\left(c_{X}c_{Y}-G_{XY}(R)\right)$
$=\sum_{X}\sum_{Y}V_{XY}(r)\left(c_{X}c_{Y}-\frac{c_{X}c_{Y}g_{XY}(R)}{\sum_{A}\sum_{B}c_{A}c_{B}g_{AB}(R)}\right)$
$=\sum_{X}\sum_{Y}V_{XY}(r)c_{X}c_{Y}\left(1-\frac{g_{XY}(R)}{\sum_{A}\sum_{B}c_{A}c_{B}g_{AB}(R)}\right)$
$=\sum_{X}\sum_{Y}V_{XY}(r)c_{X}c_{Y}\left(1-\Phi_{XY}(R)\right),$ (66)
where $\Phi_{XY}(R)$ is defined in Eq. (15). The difference in the energy per atom between the prediction of the AA method and LOAA is then
$\Delta E_{\rm atom}$ $=\frac{1}{2}\sum_{n}N_{n}\Delta V_{\rm eff}(R_{n})$
$=\frac{1}{2}\sum_{n}N_{n}\sum_{X}\sum_{Y}V_{XY}(R_{n})c_{X}c_{Y}\left(1-\Phi_{XY}(R_{n})\right),$ (67)
where $N_{n}$ is the number of atoms in neighbor shell $n$, at distance $R_{n}$ in the reference configuration from an atom (all atoms in a hexagonal lattice have the same environment, so are equivalent). The results in Fig. 9(a) are for a binary hexagonal lattice with LJ interactions up to second neighbors. To simplify the expressions, we neglect second-neighbor interactions (which are small) so that Eq. (67) is
$\Delta E_{\rm atom}=3\left[V_{AA}(a)c_{A}^{2}(1-\Phi_{AA}(a))+2V_{AB}(a)c_{A}c_{B}(1-\Phi_{AB}(a))+V_{BB}(a)c_{B}^{2}(1-\Phi_{BB}(a))\right].$ (68)
Substituting in $V_{XY}(r)$ from Eq. (34) and normalizing by $\epsilon_{AA}$, we have
$\frac{\Delta E_{\rm atom}}{\epsilon_{AA}}=12\varphi(a)\left[\eta+\gamma\frac{\epsilon_{AB}}{\epsilon_{AA}}+\zeta\frac{\epsilon_{BB}}{\epsilon_{AA}}\right].$ (69)
where
$\eta=c_{A}^{2}(1-\Phi_{AA}(a)),\qquad\gamma=2c_{A}c_{B}(1-\Phi_{AB}(a)),\qquad\zeta=c_{B}^{2}(1-\Phi_{BB}(a)).$ (70)
Note that parameters $\eta$, $\gamma$, $\zeta$ do not depend on the potential parameters ($\epsilon_{XY}$), since the same partial RDF is observed for all $\epsilon$ values within a given LO region (as explained in Section III.1.3) and hence $\Phi_{XY}(a)$ are the same as well.
Figure 22: Change of variables in $\epsilon$-space to coordinates $s$ and $t$ that are parallel and perpendicular to the no-LO line in the LO phase diagram.
We wish to study how the energy difference in Eq. (69) changes with vertical distance from the no-LO line to understand the observations in Fig. 9(a). To this end, we perform a change of variable in $\epsilon$-space from $\epsilon_{AB}/\epsilon_{AA}$ and $\epsilon_{BB}/\epsilon_{AA}$ to the coordinates $s$ and $t$ that are parallel and perpendicular to the no-LO line as shown in Fig. 22. The equation of the no-LO line is
$\frac{\epsilon_{BB}}{\epsilon_{AA}}=2\frac{\epsilon_{AB}}{\epsilon_{AA}}-1.$ (71)
The transformation from basis vectors $\bm{e}_{1}$ and $\bm{e}_{2}$ aligned with the horizontal and vertical axes in Fig. 22 to a new basis where $\bm{e}_{1}^{\prime}$ is aligned with the no-LO line is
$\begin{bmatrix}\bm{e}_{1}^{\prime}\\ \bm{e}_{2}^{\prime}\end{bmatrix}=\begin{bmatrix}\cos\theta&\sin\theta\\ -\sin\theta&\cos\theta\end{bmatrix}\begin{bmatrix}\bm{e}_{1}\\ \bm{e}_{2}\end{bmatrix}$ (72)
where $\theta=\tan^{-1}2=63.43^{\circ}$ is the angle between $\bm{e}_{1}$ and the no-LO line, for which $\cos\theta=1/\sqrt{5}$ and $\sin\theta=2/\sqrt{5}$. Referring to Fig. 22, the position of an arbitrary point P is
$P$ $=0.5\bm{e}_{1}+s\bm{e}_{1}^{\prime}+t\bm{e}_{2}^{\prime}$
$=(0.5+s\cos\theta-t\sin\theta)\bm{e}_{1}+(s\sin\theta+t\cos\theta)\bm{e}_{2}.$ (73)
We also have that
$P=\frac{\epsilon_{AB}}{\epsilon_{AA}}\bm{e}_{1}+\frac{\epsilon_{BB}}{\epsilon_{AA}}\bm{e}_{2}.$ (74)
Equating Eqns. (73) and (74), we find
$\frac{\epsilon_{AB}}{\epsilon_{AA}}=\frac{1}{2}+\frac{1}{\sqrt{5}}(s-2t),\qquad\frac{\epsilon_{BB}}{\epsilon_{AA}}=\frac{1}{\sqrt{5}}(2s+t).$ (75)
Substituting these expressions into Eq. (69) and rearranging gives
$\frac{\Delta E_{\rm atom}}{\epsilon_{AA}}=12\varphi(a)\left[\eta+\frac{1}{2}\gamma+\frac{1}{\sqrt{5}}(2\zeta+\gamma)s+\frac{1}{\sqrt{5}}(\zeta-2\gamma)t\right].$ (76)
Thus in general we expect an approximately [^15] We are only considering near-neighbor interactions and are using the LOAA solution as an approximation for the exact TS solution. linear variation of the per-atom-energy difference between the AA and exact result moving parallel and perpendicular to the no-LO line. However, Fig. 9(a) shows contour lines parallel to the no-LO lines, which implies no dependence on $s$. This requires that
$2\zeta+\gamma=0.$ (77)
Computing this value across $\epsilon$-space, it is found to be of order $10^{-3}$$10^{-4}$ in both LO regions. In contrast the term multiplying the $t$ coordinate ($\zeta-2\gamma$) is of order $0.1$. This explains the observations in Fig. 9(a).
Appendix D Elastic Constant Equations for AA and LOAA
In this section, we derive analytical expressions for the elastic constants prediction by the AA and LOAA approaches and demonstrate their linear dependence on distance from the no-LO line. We begin with the AA approach. Replacing $V_{\rm eff}$ in Eq. (46) for AA, we have
$c_{ijkl}^{\rm AA}=\frac{1}{2\Omega}\sum_{\begin{subarray}{c}\alpha,\beta\\ \alpha\neq\beta\end{subarray}}Q^{\alpha\beta}_{ijkl}\left[\sum_{X}\sum_{Y}\left(V_{XY}^{\prime\prime}(r^{\alpha\mathring{\beta}})-\frac{V_{XY}^{\prime}(r^{\alpha\mathring{\beta}})}{r^{\alpha\mathring{\beta}}}\right)c_{x}c_{y}\right],$ (78)
where for simplicity we denote,
$Q^{\alpha\beta}_{ijkl}=\frac{r_{i}^{\alpha\mathring{\beta}}r_{j}^{\alpha\mathring{\beta}}r_{k}^{\alpha\mathring{\beta}}r_{l}^{\alpha\mathring{\beta}}}{(r^{\alpha\mathring{\beta}})^{2}}.$ (79)
Simplifying further by expanding and pulling terms out of the summations, we have
$c_{ijkl}^{\rm AA}=\frac{2}{\Omega}\left[c_{A}^{2}\epsilon_{AA}+2c_{A}c_{B}\epsilon_{AB}+c_{B}^{2}\epsilon_{BB}\right]\sum_{\begin{subarray}{c}\alpha,\beta\\ \alpha\neq\beta\end{subarray}}Q^{\alpha\beta}_{ijkl}\left[\varphi^{\prime\prime}(r^{\alpha\mathring{\beta}})-\frac{\varphi^{\prime}(r^{\alpha\mathring{\beta}})}{r^{\alpha\mathring{\beta}}}\right].$ (80)
We rewrite this as
$\frac{c_{ijkl}^{\rm AA}}{\epsilon_{AA}}=\frac{2}{\Omega}\left[\eta^{\rm AA}+\gamma^{\rm AA}\frac{\epsilon_{AB}}{\epsilon_{AA}}+\zeta^{\rm AA}\frac{\epsilon_{BB}}{\epsilon_{AA}}\right]\sum_{\begin{subarray}{c}\alpha,\beta\\ \alpha\neq\beta\end{subarray}}Q^{\alpha\beta}_{ijkl}\left[\varphi^{\prime\prime}(r^{\alpha\mathring{\beta}})-\frac{\varphi^{\prime}(r^{\alpha\mathring{\beta}})}{r^{\alpha\mathring{\beta}}}\right],$ (81)
where
$\eta^{\rm AA}=c_{A}^{2},\qquad\gamma^{\rm AA}=2c_{A}c_{B},\qquad\zeta^{\rm AA}=c_{B}^{2}.$ (82)
The motivation for this substitution will be more evident in the LOAA formulation below. Following the same change of variables procedure in Appendix C, this can be re-written as
$\frac{c_{ijkl}^{\rm AA}}{\epsilon_{AA}}=\frac{2}{\Omega}\left[\eta^{\rm AA}+\frac{1}{2}\gamma^{\rm AA}+\frac{1}{\sqrt{5}}\left(\gamma^{\rm AA}+2\zeta^{\rm AA}\right)s+\frac{1}{\sqrt{5}}\left(\zeta^{\rm AA}-2\gamma^{\rm AA}\right)t\right]\sum_{\begin{subarray}{c}\alpha,\beta\\ \alpha\neq\beta\end{subarray}}Q^{\alpha\beta}_{ijkl}\left[\varphi^{\prime\prime}(r^{\alpha\mathring{\beta}})-\frac{\varphi^{\prime}(r^{\alpha\mathring{\beta}})}{r^{\alpha\mathring{\beta}}}\right],$ (83)
where $s$ and $t$ are distances parallel and perpendicular to the no-LO line as shown in Fig. 22. A similar analysis for LOAA yields,
$\frac{c_{ijkl}^{\rm LOAA}}{\epsilon_{AA}}=\frac{2}{\Omega}\sum_{\begin{subarray}{c}\alpha,\beta\\ \alpha\neq\beta\end{subarray}}Q^{\alpha\beta}_{ijkl}\left(\eta^{\rm LOAA}+\gamma^{\rm LOAA}\frac{\epsilon_{AB}}{\epsilon_{AA}}+\zeta^{\rm LOAA}\frac{\epsilon_{BB}}{\epsilon_{AA}}\right)\left[\varphi^{\prime\prime}(r^{\alpha\mathring{\beta}})-\frac{\varphi^{\prime}(r^{\alpha\mathring{\beta}})}{r^{\alpha\mathring{\beta}}}\right],$ (84)
where
$\eta^{\rm LOAA}=c_{A}^{2}\Phi_{AA}(r^{\alpha\mathring{\beta}}),\qquad\gamma^{\rm LOAA}=2c_{A}c_{B}\Phi_{AB}(r^{\alpha\mathring{\beta}}),\qquad\zeta^{\rm LOAA}=c_{B}^{2}\Phi_{BB}(r^{\alpha\mathring{\beta}}).$ (85)
Then after the change of variables to $s$ and $t$ coordinates,
$\frac{c_{ijkl}^{\rm LOAA}}{\epsilon_{AA}}=\frac{2}{\Omega}\sum_{\begin{subarray}{c}\alpha,\beta\\ \alpha\neq\beta\end{subarray}}Q^{\alpha\beta}_{ijkl}\left[\varphi^{\prime\prime}(r^{\alpha\mathring{\beta}})-\frac{\varphi^{\prime}(r^{\alpha\mathring{\beta}})}{r^{\alpha\mathring{\beta}}}\right]\times\\ \left(\eta^{\rm LOAA}+\frac{1}{2}\gamma^{\rm LOAA}+\frac{1}{\sqrt{5}}\left(\gamma^{\rm LOAA}+2\zeta^{\rm LOAA}\right)s+\frac{1}{\sqrt{5}}\left(\zeta^{\rm LOAA}-2\gamma^{\rm LOAA}\right)t\right).$ (86)